# Operationalizing Contextual Integrity in Privacy-Conscious Assistants

### v1

@@ 4.2 Synthetic Form Generation
To generate synthetic online forms we first compiled $14$ hypothetical scenarios like "Apply for medical school", "In-person work event registration", "Create checking account in the US bank", or "Newsletter subscription to online clothes shopping website" that may plausibly relate online form filling tasks. Given one of these scenarios, we prompt Gemini Ultra to generate first a more detailed description for the form and then a title; this process is repeated $3$ times with independent random seeds for each scenario. To further increase the number of forms and enable us to test the assistant’s robustness to different phrasings, we also ask the model to produce $2$ alternative paraphrasings for each title-description pair, resulting in a total of $126$ form settings.
The fields that can appear in forms are obtained from the list of keys of information available to every persona. For each field we prompt Gemini Ultra to generate at most $5$ different ways of asking for that information in a form, e.g. date_of_birth gets mapped to "Date of birth", "Birthday", "D.O.B.". These phrasings are obtained independently from the form application and re-used across multiple forms (see Table C.1 for a full list); this step ensures the filling task is more complex than a simple pattern matching between keys. Generated field phrasings were manually reviewed to remove ambiguous or confusing phrasings, resulting in a total of $193$ phrasings.
Each form in the dataset is obtained by using one of the possible title-description pairs and generating a form by randomly selecting $7$ relevant keys and then assigning a random phrasing to each key. We keep the same keys but resample the phrasings used to ask for the key across form title and description paraphrasings. Relevance of keys for the different form scenarios is obtained via human annotations (cf. Section 4.3). To obtain the filling ground truth we retain the names of the information keys used to construct the form, and map these to ground truth values for every field in the form given a concrete persona. Note that the information of which fields correspond to which key is only used for evaluation and not made available to the assistant.
@@ 4.3 Human Annotations
Determining the appropriateness of sharing the value of a particular user information field in a certain form is a challenging problem. In our benchmark we compile ground truth labels by relying on annotations from $8$ human raters. Each rater is asked to provide labels by judging all possible $51\times 14=714$ pairs of information keys and form scenarios on a five point Likert scale. To guide raters towards judgements that represent expectations across the population (instead of individual preferences) we ask for two types of labels: necessary, meaning the field is not filled the purpose of the form cannot be achieved; and relevant, meaning the field occurs in at least some forms for the given application, although it could be optional. See Appendix C.4 for the full instructions.
@@
Deciding whether information types are necessary or relevant depends on two main factors, the context (i.e. the form, and social norms or regulations), and the user’s expectations. The second factor suggests that annotators might differ in their judgements even when the labelling task is designed to elicit norms that are valid across a certain population. This posed a challenge in extracting ground truth labels from a set of annotations - we refer to Appendix C.4 and Figure 12 for an analysis of annotator disagreements across different types of information. For the results presented in the main part of the paper, we assign labels: Yes for a data key and form application pair if all raters assign a value of 3 or higher; No if all raters assign a value of 3 or smaller; and Unsure otherwise. For “necessary” annotations this yields 50 Yes, 460 No and 232 Unsure.
@@ 5 Experimental Evaluation
We implement form filling assistants by prompting Gemini Pro and Ultra geminiteam2023gemini to play the roles of assistant, supervisor and filler described in Section 3. In particular, we evaluate the four assistant architectures we propose, including two versions of the assistant with reasoning supervisor (with and without chain-of-thought reasoning kojima2022large). We experiment with prompt engineering, few shot prompting, and the inclusion of user information in the prompt of the supervisor. In all our implementations the form-filling assistants process each field independently of the others, and supervisors can output three decisions: fill, do not fill, or ask the user - the latter expresses uncertainty in the model’s decision and is considered correct when the ground truth label is Unsure. Full experimental details (including prompts) are provided in Appendix A. Unless we say otherwise, all experiments report results over three random seeds.
@@
Evaluating the different assistant architectures using Gemini Ultra shows that all assistants except the self-censoring one achieve strong utility and privacy performance, with the CI-based reasoning supervisor outperforming the others on both utility and privacy by a small margin (Figure 4). We also examine how privacy and utility trade-off based on instructions provided in the prompt (Figure  10) and across the number of shots used in few-shot prompting (Figure 11). When using the smaller Gemini Pro model we observe a similar qualitative behavior (Table 1). There is no clear winner–rather, different assistants display different strengths, e.g. the reasoning supervisor is best in terms of privacy leakage, while the CI-based supervisor is best in terms of utility. As a baseline we also implement a form-filling assistant without any privacy mitigation using the same filler prompts and models. This baseline fills 93.7% of all the fields correctly, showing that there is room for improvement in the utility of our assistants.
Figure 3: Example of our form filling query from our benchmark and how different assistants respond. We notice that in some cases (as here), the reasoning supervisor fails to provide a reasoning and only replies with its decision. Even with CoT the model output does not add any additional explanations. The CI-based supervisor provides more structured reasoning that can, for example, be used to interpret model failures.
@@
To evaluate robustness of different assistants we investigate how their privacy judgements change across paraphrasings of the same form in Figure 5, where we observe that the CI-based supervisor assistant’s decisions exhibit less variation across paraphrasings than other assistants. Next we analyze robustness of assistant privacy judgements to changing the underlying model between Gemini Ultra and Gemini Pro and observe that the assistants with reasoning-based supervisor produce outputs that are more consistent across model size (Figure 7, left). In addition, Figure 7 (right) shows rates of agreement between different instantiations of the assistant, with agreement being overall larger between reasoning-based assistants.
Table 1: Mean and standard error of privacy leakage and utility of assistants for Gemini Pro and Ultra models. Pareto-optimal assistants per model are bold. We observe that assistants based on Gemini Ultra perform better than those based on Gemini Pro. Further reasoning supervision consistently helps in decreasing privacy leakage.
@@
The disagreement between raters expresses ambiguity or dispute about the necessity of exposing a specific piece of data in a given context. A desirable property of a universal assistant would be to exhibit higher uncertainty for information flows in dispute. As shown in Figure 6, we see that the Ultra models’ uncertainty is positively correlated with raters’ disagreement as opposed to their Pro models counterparts for all prompting approaches besides the CI approach.
@@ 6 Conclusion
AI assistants that can undertake information-sharing tasks on behalf of users can provide significant value in innumerable applications. Our work makes progress towards information-sharing assistants whose actions align with users’ privacy expectations by 1) identifying contextual integrity as an appropriate framework to ground such alignment on societal information norms, and 2) proposing a method to operationalize CI-based reasoning in LLM-based assistants. Our experimental evaluation shows that assistants built on existing models can readily benefit from this approach to significantly reduce privacy leakage without major deteriorations in utility, and in particular that CI-based assistants achieve the best performance among the options we investigate. We conclude by identifying a number of research challenges that need to be addressed by future work, and note that our work only touches upon one of many important aspects that need to be considered when responsibly developing and deploying advanced AI assistants to have positive broader impact gabriel2024ethics.
@@ 6.1 Limitations and Future Work
Improving and benchmarking CI capabilities requires high-quality data sets. We demonstrate that a combination of synthetic data and human annotations might offer a path towards larger benchmarks covering more application domains and information norms. Ideally, data sets constructed in future work will enable assistants to generalize across contexts, have a more robust notion of uncertainty (e.g. to ask for user intervention in uncertain cases), and extend CI capabilities beyond the form-filling application we consider. Norm elucidation is another critical component of benchmark construction which can benefit from future research in multiple directions, including case-by-case human annotations providing examples of norm application, as well as, explicitly documented norms that can be directly consumed by assistants (e.g. as instructions in the prompt).
@@
The assistants we propose in Section 2 can be developed via e.g. prompt engineering, few-shot prompting, or fine-tuning. Our evaluation is limited to prompt engineering and few-shot prompting as these are effective techniques that work well in the low data regime we find ourselves in. Investigating the most effective way to combine high quality data with fine-tuning-based techniques to deliver improvements in these capabilities is an important direction for future research. Another important direction for future work is to harden our designs to still protect user information when faced with a malicious third party attempting to exploit known LLM vulnerabilities – this could be achieved using the techniques from bagdasaryan2024air.
@@
CI identifies the features of an information flow that are sufficient to determine whether the appropriate norms are being followed. However, these features can be as rich and complex as the range of applications covered by the assistant, making it difficult to compile a complete taxonomy for the values that features like “receiver”, “information type” or “context” can take in the IFC. This poses important challenges for future work on data collection and evaluation methodologies of CI capabilities in AI assistants, especially in contrast with works that consider narrower applications where taxonomies of CI-relevant features can be manually constructed DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq; DBLP:conf/chi/AbdiZRS21. Finally, while CI only concerns itself with widely accepted norms, personalization is a critical aspect of assistive technologies whose operationalization might require going beyond CI but could still benefit from our information-sharing assistant abstraction.
@@
We thank Taylan Cemgil, Johannes Welbl, Leo Cheng and many others for their valuable inputs into this project.
@@ Appendix B Additional Experimental Results
We provide an overview of all results obtained through changes in prompting in Table 5. In particular, Figures 9, 10 and 11 visualise how changes in the user information, the prompting and the few shot examples influence the assistants’ performance.
@@
Figure 10: In the results of the main paper, the assistants are prompted to decide on the necessity of a data key. Here we provide the models with the instructions that were shared with the human instructors (see Figures 13 and 14). We observe that privacy leakage decreases at the cost of decreased utility.
@@ B.1 Adjusting to User Preferences
Figure 12: Ratio of “necessary” to “relevant” labels across different forms per each rater. Rater 1 has a lower share of “necessary” fields compared to other raters. Form 7 for newsletter shopping subscription has the lowest share for necessary over relevant fields.
@@ C.2 Form Generation
     Paraphrase the following text without changing its meaning. Replace words’  with synonymes, change proprietary names, add words that do not change the meaning, change the order of words, change nouns to verbs or viceversa, change capitalisation, or include additional punctuation, spelling or grammar mistakes.    Examples:  text: Job Application as Apprentice at Virgin Media.  paraphrased text: Apply for a job as Apprentice at Sky.    text: Complaint Form for our Customers.  paraphrased text: Clients’ Complaint Form.    text: Registration for Virtual AI workshop.  paraphrased text: Sign up for Online ML Workshop.    text: Contact Support.  paraphrased text: Contact the support team.    text: {TITLE}  paraphrased text:
@@ C.4 Human Annotations
We collected 8 sets of human ratings for each combination of webform application and data key. We collected two sets of labels from each human rater: necessary and relevant. Please check Figures 13 and 14 for the instructions as provided to the human raters.

### v2

@@ 4.3 Synthetic Form Generation
To generate synthetic online forms we first compiled $14$ scenarios like "In-person work event registration", "Create checking account in the US bank", or "Newsletter subscription to online clothes shopping website" that mimic real online form filling tasks. Given one of these scenarios, we prompt Gemini Ultra to generate first a more detailed description for the form and then a title; this process is repeated with $3$ random seeds for each scenario. To further increase the number of forms and enable us to test the assistant’s robustness to different phrasings, we also ask the model to produce $2$ alternative paraphrasings for each title-description pair, resulting in $126$ forms.
The fields that can appear in forms are obtained from the list of keys of information available to every persona. For each field we prompt Gemini Ultra to generate $5$ different ways of asking for that information in a form, e.g. date_of_birth gets mapped to "Date of birth", "Birthday", "D.O.B.". These phrasings are obtained independently from the form application and re-used across multiple forms (see Table C.1 for a full list); this step ensures the filling task is more complex than a simple pattern matching between keys. Generated field phrasings were manually reviewed to remove ambiguous or unrealistic phrasings, resulting in $193$ phrasings.
Each form in the dataset is obtained by using one of the possible title-description pairs and generating a form by randomly selecting $7$ relevant keys and then assigning a random phrasing to each key. To obtain the filling ground truth we retain the names of the information keys used to construct the form, and map the to ground truth values for every field in the form given a concrete persona. We keep the same keys but resample the way fields are phrased in the form together with each form title and description paraphrasings. Note that the information of which fields correspond to which key is only used for evaluation and not made available to the assistant. Relevance of keys for the different form scenarios is obtained via human annotations (cf. Section 4.4).
@@ 4.4 Human Annotations
Determining the appropriateness of sharing the value of a particular user information field in a certain form is a challenging problem. In our benchmark we compile ground truth labels by relying on annotations from $8$ human raters. Each rater is asked to provide labels by judging all possible $51\times 14=714$ pairs of information keys and form scenarios on a five point Likert scale. To guide raters towards judgements that represent expectations across the population (instead of individual preferences) we ask for two types of labels: necessary, meaning the field is not filled the purpose of the form cannot be achieved; and relevant, meaning the field occurs in at least some forms for the given application, although it could be optional.
Figure 4: Ratio of “necessary” to “relevant” labels across different form applications F1-F14 per each rater. The form applications can be found in Supplementary Table 7.
@@
Deciding whether information types are necessary or relevant depends on two main factors, the context (i.e. the form, and social norms or regulations), and the user’s expectations. The second factor suggests that annotators might differ in their judgements even when the labelling task is designed to elicit norms that are valid across a certain population (see Figure 4). This posed a challenge in extracting ground truth labels from a set of annotations - we refer to Appendix C.4 for an analysis of annotations across different aggregation strategies and more analysis on raters’ disagreement. As we try to capture shared social norms, we deem a data key as necessary only if none of the raters disagrees with its necessity. For “necessary” annotations this yields 50 Yes, 460 No and 232 Unsure.
@@ 5 Quantitative Analysis
We implement form filling assistants by prompting Gemini geminiteam2023gemini, Gemma 2 (team2024gemma) and Mistral models (jiang2023mistral), in particular mistral-small-2402 and mistral-large-2402, to play the roles of assistant, supervisor and filler described in Section 3. In particular, we evaluate the four assistant architectures we propose, including two versions of the assistant with reasoning supervisor (with and without chain-of-thought kojima2022large). We experiment with prompt engineering, few shot prompting, and the inclusion of user information in the prompt of the supervisor. In all our implementations the form-filling assistants process each field independently of the others, and supervisors can output three decisions: fill, do not fill, or ask the user - the latter expresses uncertainty in the model’s decision and is considered correct when the ground truth label is Unsure. Full experimental details (including prompts) are provided in Appendix A. Unless we say otherwise, all experiments report results averaged over three random seeds.
@@
Evaluating the different assistant architectures shows that all assistants except the self-censoring one achieve strong utility and privacy performance, with the CI-based reasoning supervisor outperforming the others on both utility and privacy by a small margin using Gemini Ultra (Figure 6). We also examine how privacy and utility trade-off based on instructions provided in the prompt (Figure  12) and across the number of shots used in few-shot prompting (Figure 13). There is no clear winner–rather, different assistants display different strengths, e.g. the reasoning supervisor is best in terms of privacy leakage, while the CI-based supervisor is best in terms of utility. When using the smaller Gemini Pro, Gemma or Mistral models we observe a similar qualitative behavior (Table 1), although the overall performance becomes slightly worse. As a baseline we also implement a form-filling assistant without any privacy mitigation using the same filler prompts and models. This Gemini Ultra baseline fills 93.7% of all the fields correctly, showing that there is room for improvement in the utility of our assistants.
Figure 5: Example of our form filling query from our benchmark and how different assistants based on Gemini Ultra respond. We notice that in some cases (as here), the reasoning supervisor fails to provide a reasoning and only replies with its decision. Even with CoT the model output does not add any additional explanations. The CI-based supervisor provides more structured reasoning that can, for example, be used to interpret model failures.
@@
To evaluate robustness of different assistants we investigate how their privacy judgements change across paraphrasings of the same form (Figure 7), where we observe that the CI-based supervisor assistant’s decisions exhibit less variation across paraphrasings than other assistants. Next we analyze robustness of assistant privacy judgements to changing the underlying model between Gemini Ultra and Gemini Pro and observe that the assistants with reasoning-based supervisor produce outputs that are more consistent across model size (Figure 9, left). In addition, Figure 9 (right) shows rates of agreement between different instantiations of the assistant, with agreement being overall larger between reasoning-based assistants.
Table 1: Mean and standard error of privacy leakage and utility of assistants for Gemini, Gemma and Mistral models. Pareto-optimal assistants per model are bold. We observe that assistants based on larger models perform better than those based on smaller models (except for Gemma). Further reasoning supervision consistently helps in decreasing privacy leakage.
@@
CI-based supervisor-CoT 0.04±0.02 0.73±0.03 0.01±0.01 0.87±0.02
Gemma 2.0 9B Gemma 2.0 27B
self-censoring 1.00±0.00 0.83±0.03 1.00±0.00 0.90±0.02
binary supervisor 0.00±0.00 0.55±0.04 0.00±0.00 0.66±0.04
reasoning supervisor 0.00±0.00 0.58±0.04 0.00±0.00 0.57±0.04
reasoning supervisor-CoT 0.02±0.01 0.69±0.03 0.00±0.00 0.59±0.03
CI-based supervisor 0.01±0.01 0.68±0.04 0.00±0.00 0.66±0.04
CI-based supervisor-CoT 0.02±0.01 0.71±0.03 0.00±0.00 0.68±0.03
Mistral 1.0 Small Mistral 1.0 Large
self-censoring 0.77±0.03 0.72±0.04 0.31±0.04 0.88±0.02
binary supervisor 0.38±0.04 0.84±0.03 0.01±0.01 0.76±0.03
reasoning supervisor 0.27±0.03 0.79±0.03 0.00±0.00 0.73±0.03
reasoning supervisor-CoT 0.26±0.03 0.80±0.03 0.02±0.01 0.81±0.03
CI-based supervisor 0.06±0.01 0.81±0.03 0.01±0.00 0.84±0.03
CI-based supervisor-CoT 0.04±0.01 0.83±0.03 0.02±0.01 0.86±0.03
@@
The disagreement between raters expresses ambiguity or dispute about the necessity of exposing a specific piece of data in a given context. A desirable property of a universal assistant would be to exhibit higher uncertainty for information flows in dispute. As shown in Figure 8, we see that the Ultra models’ uncertainty is positively correlated with raters’ disagreement as opposed to the Pro models for all prompting approaches besides the CI approach.
@@ 6 Conclusion
AI assistants that can undertake information-sharing tasks on behalf of users can provide significant value in innumerable applications. Our work makes progress towards information-sharing assistants whose actions align with users’ privacy expectations by 1) identifying CI as an appropriate framework to ground such alignment on societal information norms, and 2) proposing a method to operationalize CI-based reasoning in LLM-based assistants. Our evaluation is a first step towards showing that assistants built on existing models can readily benefit from this approach to significantly reduce privacy leakage without major deterioration in utility, and in particular that CI-based assistants achieve the best performance among the options we investigate. Our work makes a first inroad to adding real world complexity into evaluation of information-sharing assistants by introducing different use cases and synthetic personas to model human interaction. This is an important step towards leveraging CI to improve alignment of human-AI interactions with user privacy expectations.
@@ 6.1 Limitations and Future Work
Improving and benchmarking CI capabilities requires high-quality data sets. We demonstrate that human annotations of synthesised data might offer a path towards larger benchmarks covering more application domains and information norms. Ideally, data sets constructed in future work will enable assistants to generalize across contexts, have a more robust notion of uncertainty (e.g. to ask for user intervention in uncertain cases), and extend CI capabilities beyond the form-filling application we consider. Norm elucidation is another critical component of benchmark construction which can benefit from future research in multiple directions, including case-by-case human annotations providing examples of norm application, as well as, explicitly documented norms that can be directly consumed by assistants (e.g. as instructions in the prompt). We rely on synthetic data and note that this is a first step for introducing human interaction to privacy evaluation, and that a next step will be to validate these findings with further experiments with humans. Future work will explore how users perceive the benefits of a CI-based form-filling assistant when filling in forms in the web.
@@
The assistants we propose in Section 2 can be developed via e.g. prompt engineering, few-shot prompting, or fine-tuning. Our evaluation is limited to prompt engineering and few-shot prompting as these are effective techniques that work well in the low data regime we find ourselves in. Investigating the most effective way to combine high quality data with fine-tuning-based techniques to deliver improvements in these capabilities is an important direction for future research. Another important direction for future work is to harden our designs to still protect user information when faced with a malicious third party attempting to exploit known LLM vulnerabilities – this could be achieved using the techniques from bagdasaryan2024air. Lastly, future research should focus on designing the user interface between the assistants and the users.
@@
CI identifies the features of an information flow that are sufficient to determine whether the appropriate norms are being followed. However, these features can be as rich and complex as the range of applications covered by the assistant, making it difficult to compile a complete taxonomy for the values that features like “receiver”, “information type” or “context” can take in the IFC. This poses important challenges for future work on data collection and evaluation methodologies of CI capabilities in AI assistants, especially in contrast with works that consider narrower applications where taxonomies of CI-relevant features can be manually constructed DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq; DBLP:conf/chi/AbdiZRS21. Finally, while CI only concerns itself with widely accepted norms, personalization and adaptation to cultural norms kiehne2022contextualizing is a critical aspect of assistive technologies whose operationalization might require going beyond CI but could still benefit from our information-sharing assistant abstraction.
@@
We thank Taylan Cemgil, Johannes Welbl, Leo Cheng, Madiha Zahrah Choksi, Daniel Votipka and many others for their valuable inputs into this project.
@@ Appendix B Additional Experimental Results
We provide an overview of all results obtained through changes in prompting in Table 5. In particular, Figures 11, 12 and 13 visualise how changes in the user information, the prompting and the few shot examples influence the assistants’ performance.
@@
Figure 12: In the results of the main paper, the assistants are prompted to decide on the necessity of a data key. Here we provide the models with the instructions that were shared with the human instructors (see Figures 14 and 15). We observe that privacy leakage decreases at the cost of decreased utility.
@@ C.2 Form Generation
      Paraphrase the following text without changing its meaning. Replace words’    with synonymes, change proprietary names, add words that do not change the meaning, change the order of words, change nouns to verbs or vice-versa, change capitalisation, or include additional punctuation, spelling or grammar mistakes.      Examples:    text: Job Application as Apprentice at Virgin Media.    paraphrased text: Apply for a job as Apprentice at Sky.      text: Complaint Form for our Customers.    paraphrased text: Clients’ Complaint Form.      text: Registration for Virtual AI workshop.    paraphrased text: Sign up for Online ML Workshop.      text: Contact Support.    paraphrased text: Contact the support team.      text: {TITLE}    paraphrased text:
@@ C.4 Human Annotations
We collected 8 sets of human ratings for each combination of webform application and data key. We collected two sets of labels from each human rater: necessary and relevant. Please check Figures 14 and 15 for the instructions as provided to the human raters.